Session: Feb 2012 (session 1)
Wavelength: 3.6 cm
To give feedback on this experiment, click
(Hua Zhang at Sun Feb 26 10:12:14 2012) Success:
Yebes (Angel Vicente at Sun Feb 26 15:03:34 2012) Minor failures: HIGH
SCAN 1 - 140 aprox. with problems Tsys
(Geert Kuper at Sun Feb 26 15:20:32 2012) Minor failures: RECO
Diskpack OSOD-046 sometimes error messages during scancheck.
Also LED o/1 was not lit during recording. Therefore switched to
fresh diskpack at 09:58 UT.
(A.Maccaferri at Sun Feb 26 15:21:48 2012) Success:
(Michael Lindqvist at Mon Feb 27 06:44:44 2012) Success:
Some bad scan_check, otherwise no known problems.
Yebes (Pablo de Vicente at Mon Feb 27 07:48:25 2012) Minor failures:RECE HIGH FS
- No IF power between 04:30 and 07:10. This part can be disregarded.
- Power was correct from 07:10
- Observation was halted between 8:54 to 9:33
- Tsys was wrongly computed before 9:33 because the noise diode attenuation was incorrectly set. It can be correctly recomputed using different caltemps.
- From 9:33 to 15:30 the observation worked correctly
(Uwe Bach at Mon Feb 27 09:18:16 2012) Success:
Badary (A.Mikhailov at Mon Feb 27 12:33:50 2012) Success:
Earthquake at 06:20 UT
Svetloe (T.Andreeva at Mon Feb 27 12:39:46 2012) Success:
Scans no0235 - no0245: mark5b formatter time error
(Bo Xia at Tue Feb 28 23:10:55 2012) Success:
Hartebeesthoek (Jonathan Quick at Thu Mar 01 12:42:34 2012) Minor failures:RECE
Some IFD communication problems. VC1 was incorrectly patched low instead of high by mistake.